What is internal resistance of cell or bettry
1
Expert's answer
2020-10-07T08:59:06-0400

Internal resistance means the opposition to the flow of current offered by the cells and batteries themselves resulting in the generation of heat.

The relationship between internal resistance (r) and emf (e) of cell s given by.

=> "e= I(r+R)"

Where, e = EMF i.e. electromotive force (Volts), I = current (A), R = Load resistance, and r is the internal resistance of cell measured in ohms.

On rearranging the above equation we get;

"=> e= IR+Ir => e= V+Ir"

In the above equation, V is the potential difference (terminal) across the cell when the current (I) is flowing through the circuit.

Note: The emf (e) of a cell is always greater than the potential difference (terminal) across the cell